OpenStudy (anonymous):

Mary's age is 2/3 that of Peter's.Two years ago Mary's age was 1/2 of what peter's age will be in 5 year's time.How old is Peter now?

OpenStudy (anonymous):

it is simple algebra~~ just let Mary's age as M and Peter's age as P. Then you will get 2 equations. First, M = (2/3) P and (M-2) = (1/2) (P+5). Then you will get that actually Peter is 27 years old now by using substitution or elimination~ hope you get it!